What is the specific heat of aluminum if the temperature of an 85.2g sample of aluminum is increased by 24.3º, when 621J of heat is added?


1
Expert's answer
2021-04-02T05:14:40-0400

Solution:

q = m × C × ΔT

where:

q = amount of heat energy gained or lost by substance (J)

m = mass of sample (g)

C = specific heat capacity (J oC-1 g-1)

ΔT = change in temperature (oC)

 

Thus:

(621 J) = (85.2 g) × C × (24.3º)

С = (621 J) / (85.2 g × 24.3oC) = 0.2999 J oC-1 g-1 = 0.300 J oC-1 g-1

C(Al) = 0.300 J oC-1 g-1


NB! Although the specific heat of aluminum should be 0.900 J oC-1 g-1 (table value).


Answer: The specific heat of aluminum is 0.300 J oC-1 g-1